In hockey, a team recieves 2 points for a win and 1 point for a tie. During a hockey season of 60 games, the Leafs lost 28 games but earned 51 points. How many games did they win? Please and thank you :)
Answer by richwmiller(17219)

2w+1t=51
w+t=60-28=32
2w+1t=51
w+t=32
subtract
w=19
t=13
l=28